### v1

@@ Abstract
Vision Transformer (ViT) has prevailed in computer vision tasks due to its strong long-range dependency modelling ability. However, its large model size with high computational cost and weak local feature modeling ability hinder its application in real scenarios. To balance computation efficiency and performance, we propose SAEViT (Sparse-Attention-Efficient-ViT), a lightweight ViT based model with convolution blocks, in this paper to achieve efficient downstream vision tasks. Specifically, SAEViT introduces a Sparsely Aggregated Attention (SAA) module that performs adaptive sparse sampling based on image redundancy and recovers the feature map via deconvolution operation, which significantly reduces the computational complexity of attention operations. In addition, a Channel-Interactive Feed-Forward Network (CIFFN) layer is developed to enhance inter-channel information exchange through feature decomposition and redistribution, mitigating redundancy in traditional feed-forward networks (FNN). Finally, a hierarchical pyramid structure with embedded depth-wise separable convolutional blocks (DWSConv) is devised to further strengthen convolutional features. Extensive experiments on mainstream datasets show that SAEViT achieves Top-1 accuracies of 76.3% and 79.6% on the ImageNet-1K classification task with only 0.8 GFLOPs and 1.3 GFLOPs, respectively, demonstrating a lightweight solution for various fundamental vision tasks.
@@ 1 Introduction
ViT has achieved remarkable results in various vision tasks (e.g. classification[31], detection[12] and segmentation[15] etc.) due to its powerful global modeling capabilities. To integrate convolution neural network (CNN) with prior knowledge and local feature extraction capability into ViT, existing methods[6] have explored several fusion schemes with great success. MetaFormer[35] proved that the architecture of ViT is beneficial for improving the performance. To enhance model accuracy, some works[25] tried to refine the original attention computation by introducing multi-scale self-attention mechanisms (e.g. integrating features of different resolutions during attention operation). However, they usually add 2–4 additional multi-scale computation modules, which causes significant computational overhead and also impedes their applications in real-time scenarios. To mitigate computation cost, traditional methods employed multi-level hierarchical architectures that down-sample feature maps progressively or adopted sliding window strategies in spatial dimension. However, they relied on a fixed sliding window, which can easily ignore the correlation of local information. Other methods[32] performed sub-sampling on keys (K) and values (V) in attention computation to further reduce the computation cost.
On the other hand, the classical feed-forward network (FFN) in ViT typically consists of fully connected layers for channel-wise computation, which lacks explicit inter-channel modeling ability. While some works[33] attempted to enhance the expressivity of FFN by appending convolutional blocks or DWSConv to improve local feature modeling, which still overlooked inter-channel interactions.
@@
To address the above mentioned challenges and design an advanced backbone for vision tasks, we propose a simple yet effective Transformer block that combines local perception and long-range modeling on the basis of multi-level pyramid structure. Firstly, we integrate convolutional feature processing modules into the model to incorporate convolutional prior knowledge. Then, we adopt a sparse attention approach for feature extraction to reduce the computational cost of attention operation. By leveraging the redundancy of similar pixels in the input image, the aggregation of local features and sparse sampling are performed to ensure long-range dependencies of attention. To mitigate potential accuracy loss by sparse attention and enhance inter-channel interactions, we replace FFN layer by CIFFN, which decomposes feature channels and reassign feature weights explicitly to preserve representational capacity under lightweight constraints.
@@
Built upon the above design principle, we propose SAEViT, a lightweight model that combines the strengths of CNNs and ViT. SAEViT originates from convolution+Trans-former hybrid structure and sparse attention, while we focus on lightweight design for edge devices by striking a balance between performance and efficiency. Specifically, SAEViT’s sparse attention module achieves faster computation than Spatial Reduction Attention (SRA) and Window-based Attention, without introducing additional computational overhead. We evaluate model performance primarily through accuracy and throughput. Throughput measures inference efficiency, reflecting the real-time applicability and computational resource utilization. Higher throughput indicates more images/data samples processed per second. We develop 2 model variants with different computational complexities: Tiny (T) and Extra Small (XS), but both have $\leq 10$M parameters and $\leq 1.5$GFLOPs. We compare our model with other lightweight models in this paper, and our model indeed has less parameters and lower computation cost than other competing methods, which is supposed to be deployed under resource restricted situations. Experiments include image classification, object detection, and semantic segmentation. On ImageNet, SAEViT-t achieves 76.3% Top-1 accuracy with 0.8 GFLOPs, while the SAEViT-xs variant reaches 79.6% with 1.3 GFLOPs. Comparison of throughput will be reported in Section 4.4.
@@
  • A unified ViT framework has been proposed, which integrates CNNs and ViT to acquire the feature prior capabilities of CNN as well as the long-range dependency modeling ability of attention mechanisms.
@@
  • A channel-interactive Feed-Forward Network has been developed to counteract potential accuracy degradation caused by sparse attention. Meanwhile, channel interaction has been performed to improve the expressivity of FNN the training accuracy under lightweight constraints.
@@ 2.1 Vision Transformer
The rise of attention mechanisms promoted the popularity of ViT in the field of computer vision. As a result, a series of excellent works have emerged [18, 29, 7], which significantly improves the performances of ViT based models (in terms of efficiency and accuracy). To enhance the overall feature extraction and generalization capabilities of models, many works focus on optimizing ViT’s computational paradigm, particularly through the fusion of convolution blocks and attention mechanisms. Extensive experiments demonstrate that integrating CNN with ViT is an effective way for feature extraction and improving performance gain. Other approaches [38, 32, 18] aim to improve feature utilization efficiency by facilitating cross-layer feature interactions.
@@ 2.2 The Feed-Forward Network
The FFN layer is a critical component in ViT. Its primary role is to introduce non-linear transformations and enhance the model’s representational capacity. By expanding and compressing channel dimensions in high-dimensional spaces, it enriches the feature representation of individual tokens within the architecture. It not only affects the model’s expressiveness but also ViT’s computational efficiency and practical deployment. Consequently, ensuing works attempted to enhance FFN layers through convolutional operations to strengthen local features (e.g. PVTv2[33], SRFormer[40], LightViT[10]). Others, from architectural perspective, modified the residual connections by designing inverted residual blocks to enhance cross-layer gradient propagation (e.g. CMT[6]). Partially inspired by the above works, we focus on designing learnable parameters of FFN that interact with input feature maps to dynamically redistribute channel-wise parameters. Combining with SAA, our approach effectively boosts feature extraction capabilities while maintaining computational efficiency.
@@ 2.3 Lightweight ViT
Considering that ViT suffers high computational cost, which hinders its application in resource constrained scenarios. Designing lightweight ViT is a research area of great significance. Current lightweight ViT designs mainly focus on latency reduction on real devices, optimizing GPU and CPU throughputs, and reducing the computational overhead of attention mechanisms and overall architecture. RepViT[30] replaced the Transformer’s multi-head self attention (MSA) module with lightweight convolutional blocks. It adopted a multi-branch structure during training to enrich feature representations, and merged these branches into an efficient convolutional layer for inference. Meanwhile, it adjusted the number of channels and network depth to reduce parameters, which turns out to realize effectively lower on device latency. MobileViT[21] employed a hybrid architecture with unfold–transform–fold operations, achieving higher accuracy while reducing FLOPs. EfficientViT[1] effectively boosted both accuracy and GPU throughput by revising the inverted bottleneck structure and introducing cascaded group attention.
@@ 3 Method
Our goal is to combine the strength of convolutional neural networks (CNN) with the attention mechanisms, and develop a lightweight ViT based architecture. Following the design principle of pyramid network, our model is divided into 4 stages, each of which produces feature maps of different scales to capture multi-scale image information for detection and segmentation tasks. Between each stage, Patch Embedding layers (which consists of convolution and normalization operations) are adopted to down-sample the resolution to half of its original value. In Section 4.5.4, the convolution stems are proved to be effective at boosting accuracies in multiple downstream vision tasks. Accordingly, our convolution stem is built upon 2 convolution layers (with kernel size=3, stride=1, padding=1). Meanwhile, we stack various Transformer blocks at each pyramid layer to meet diverse task requirements and computational resource constraints. Each Transformer blocks share the same internal structure, which is comprised of a SAA module and a CIFFN. The overall architecture of our proposed model is shown in Fig. 1.
@@
To construct a hybrid convolution-ViT structure, we connect Local Feature Extraction (LFE)[36] module with modified Transformer block. LFE is an embedding layer with DWSConv and an activation function, which refines the input features to the Transformer block. Inside the Transformer block, SAA module rectifies the sampling strategy by performing the sparse block-wise sampling on the feature map in an adaptive way, which reduces the computational load of traditional attention mechanisms. To deal with information redundancy and limited feature expressivity in high-dimensional features, we rectify FNN by employing inter-channel interaction mechanism. To be specific, a 1×1 convolution is firstly applied to map the input features into a higher dimensional space. Then, depth-wise convolution (DWConv) is used to enhance features within each channel, enriching the feature content. An activation function is then introduced to introduce non-linearity, further improving the model’s fitting ability. During channel interaction, a customized decomposition layer and an element-wise scaling layer are used to redistribute and adjust features across channels, enhancing inter-channel information exchange. Finally, another 1×1 convolution maps the features back to their original dimension. Our approach significantly reduces information redundancy and enhances the model’s capability to represent complex features.
The extracted features are sent to a global average pooling layer to yield global feature representation, which is then followed by a classification layer with a softmax function that ultimately produce the final classification results. Based on this, SAEViT effectively addresses the inefficient feature processing mechanism in traditional ViT, which also provides an efficient and lightweight solution for vision tasks.
@@ 3.1 Sparsely Aggregated Attention
The core idea of SAA originates from the fact that in vision-related tasks, adjacent image patches often have high redundancies (i.e. they share high similarities). To mitigate the redundancy and also to capture long-range spatial dependencies, we replace the standard attention block with SAA in our lightweight model, so that the model will be able to filter out and integrate tokens that carry similar information (i.e. reduce the number of tokens and computation costs). SAA is featured by spatial sampling strategy: by down-sampling the tokens with a rate sr, so that only the features in the key regions are retained. To be specific, the stride of down-sampling is set to sr to aggregate the adjacent tokens with high similarity. Then, each non-overlapping sr × sr window is reduced to 1 token through average pooling. It not only reduces the complexity of attention operation, but also preserves the main information within the window through local feature smoothing, avoiding the loss of key details caused by random sampling. This is similar to MaxViT[23]’s "Sparse Global Attention" and Biformer[33]’s "Dynamic Sparse Routing" approach, both of them reduce computational complexity through spatial down-sampling. However, unlike them, SAA later up-samples the feature map via deconvolution so as to recover the original spatial dimension. Deconvolution takes a small sized input feature map and outputs a larger feature map, which restores the spatial structural information through a learnable convolution kernel (i.e. enlarging the size by inserting zeros between input elements and performing standard convolution). For given sizes of input feature maps, if the stride is set to s, then the intermediate expended feature map could be created by inserting s-1 number of zeros between each row and column of each spatial unit. Then, a learnable convolution kernel with a size of k×k is used to convolve the feature map and the output size is controlled by boundary padding.
@@ 3.2 Channel Interaction Feedforward Network Layer
A standard FFN is typically composed of fully connected layers. However, redundancy exists across different channels when dealing with high-dimensional features. In addition, the linear transformation in fully connected layers applies the same operation to all channels, lacking effective screening mechanism and integration of inter-channel information. As a result, the model may overly focus on some unimportant information during training while ignoring the correlations between channels, compromising the model’s expressive power and generalization abilities. To address this issue, most existing methods directly implant channel enhancement modules, such as the Squeeze-and-Excitation (SE) module, into the FFN. Meanwhile, the non-linear transformations in conventional FFN often apply activation functions to the outputs of fully connected layers, which may only capture insufficient multi-scale and hierarchical features under complex feature distributions. Under such circumstances, we propose CIFFN, which consists of convolutional layers for channel mapping, DWConv for local information extraction, and a feature decomposition module for decomposing and recombining features. A comparison has been made among FFN, depth-wise separable (DWS) FFN and CIFFN in Fig. 3. The computation flow of CIFFN is expressed as:
@@ 4 Experiments
In this section, multiple visual recognition experiments have been conducted to verify the effectiveness of our proposed model. First, our model is pre-trained on ImageNet-1K[4] for image classification task, which is compared with other popular in terms of accuracy and computational cost. Besides, the performances for object detection and segmentation are also tested on COCO [17] and ADE20K[39] respectively. The processing speeds for feature maps are analyzed and compared with other competitors.
@@ 4.1 Image Classification
Results: A comparison of classification results are made between our model and other popular methods, including MobileNetV2[26] and EfficientNet[27] as well as PVT[32], PVTv2[33], and DeiT[28]. All methods are roughly divided into 2 groups based on their number of parameters and floating-point operations (FLOPs). As shown in Table 2, under comparable network size, SAEViT achieves higher accuracy than other peers.
@@
CaiT-xxs12-withDW[37] 6.6 1.3 75.89
Incepformer-S[22] 16.07 9.41 75.37
@@
CaiT-xxs24-withDW[37] 12.0 2.5 79.66
Incepformer-M[22] 30.96 19.2 77.58

### v2

@@ Abstract
Vision Transformer (ViT) has prevailed in computer vision tasks due to its strong long-range dependency modelling ability. However, its large model size and weak local feature modeling ability hinder its application in real scenarios. To balance computation efficiency and performance in downstream vision tasks, we propose an efficient ViT model with sparse attention (dubbed SAEViT) and convolution blocks. Specifically, a Sparsely Aggregated Attention (SAA) module has been proposed to perform adaptive sparse sampling and recover the feature map via deconvolution operation, which significantly reduces the computational complexity of attention operations. In addition, a Channel-Interactive Feed-Forward Network (CIFFN) layer is developed to enhance inter-channel information exchange through feature decomposition and redistribution, which mitigates the redundancy in traditional feed-forward networks (FFN). Finally, a hierarchical pyramid structure with embedded depth-wise separable convolutional blocks (DWSConv) is devised to further strengthen convolutional features. Extensive experiments on mainstream datasets show that SAEViT achieves Top-1 accuracies of 76.3% and 79.6% on the ImageNet-1K classification task with only 0.8 GFLOPs and 1.3 GFLOPs, respectively, demonstrating a lightweight solution for fundamental vision tasks.
@@ 1 Introduction
ViT has achieved remarkable results in various vision tasks (e.g. classification[31], detection[12] and segmentation[15] etc.) due to its powerful global modeling capabilities. Existing methods (e.g.[6], [35]) tried to integrate convolution neural network (CNN) with prior knowledge and local feature extraction capability into ViT. To enhance model accuracy, Ren et al.[25] refined the original attention computation by introducing multi-scale self-attention mechanisms. However, the additional multi-scale computation modules would cause significant computational overhead, which impede their applications in real-time scenarios. To reduce computation cost, traditional methods employed multi-level hierarchical architectures that down-sample feature maps progressively or adopted sliding window strategies in spatial dimension. However, they relied on a fixed sliding window, which can easily ignore the correlation of local information. Other methods[32] performed sub-sampling on keys (K) and values (V) in attention computation to further reduce the computation cost. On the other hand, Wang et al.[33] attempted to enhance the expressivity of FFN by appending convolutional blocks or DWSConv to improve local feature modeling, but it overlooked inter-channel interactions.
@@
To address the above mentioned problems, we propose a simple yet effective backbone for multiple vision tasks, where we endow ViT with local perception and long-range modeling ability. Firstly, we integrate convolutional feature processing modules into the model to incorporate convolutional prior knowledge. Then, we adopt a sparse attention approach for feature extraction to reduce the computational cost of attention operation. By leveraging the redundancy of similar pixels in the input image, the aggregation of local features and sparse sampling are performed to ensure long-range dependencies of attention. Finally, to alleviate potential accuracy loss caused by sparse attention and to enhance inter-channel interactions, we replace FFN layer by CIFFN, which decomposes feature channels and reassign feature weights explicitly to preserve representational capacity under lightweight constraints.
@@
Built upon the above design principle, we propose SAEViT, a lightweight model that combines the strengths of CNNs and ViT. SAEViT originates from convolution+Trans-former hybrid structure and sparse attention, while we focus on lightweight design for edge devices by striking a balance between performance and efficiency. Specifically, SAEViT’s sparse attention module achieves faster computation than Spatial Reduction Attention (SRA) and Windowbased Attention, while it does not introduce additional computational overhead. We evaluate model performance primarily through accuracy and throughput. Throughput measures inference efficiency, reflecting the real-time applicability and computational resource utilization. Higher throughput indicates more images/data samples processed per second. We develop 2 model variants with different computational complexities: Tiny (T) and Extra Small (XS), but both have $\leq 10$M parameters and $\leq 1.5$GFLOPs. Aiming at resource restricted situations, we compare our model with other lightweight models. It turns out that SAEViT achieves better performance, while it indeed has fewer number of parameters and lower computation cost. Experiments include image classification, object detection, and semantic segmentation. Take classification for instance, SAEViT-t achieves 76.3% Top-1 accuracy with 0.8 GFLOPs on ImageNet, while the SAEViT-xs variant reaches 79.6% with 1.3 GFLOPs. Comparison of throughput and latency will be reported in Section 4.4.
@@
  • A novel ViT framework has been proposed by combining CNN with ViT. It possesses the feature prior capabilities of CNN as well as the long-range dependency modeling ability of attention mechanisms.
@@
  • A channel-interactive Feed-Forward Network has been developed to counteract potential accuracy degradation caused by sparse attention. Meanwhile, channel interaction has been performed to improve the expressivity of FNN under lightweight constraints.
@@ 2.1 Vision Transformer
The rise of attention mechanisms promoted the popularity of ViT in the field of computer vision. As a result, a series of excellent works have emerged [18, 29, 7], which significantly improves the performances of ViT based models (in terms of efficiency and accuracy). To enhance the overall feature extraction and generalization capabilities of models, many works focus on optimizing ViT’s computational paradigm, particularly through the fusion of convolution blocks and attention mechanisms. Extensive experiments demonstrate that integrating CNN with ViT improves feature extraction. Other approaches [38, 32, 18] aim to improve feature utilization efficiency by facilitating cross-layer feature interactions.
@@ 2.2 The Feed-Forward Network
The FFN layer is a critical component in ViT. Its primary role is to introduce non-linear transformations and enhance the model’s representational capacity. By expanding and compressing channel dimensions in high-dimensional spaces, it enriches the feature representation of individual tokens within the architecture. It affects both the model’s expressiveness and ViT’s computational efficiency. Subsequent works attempted to enhance FFN layers through convolutional operations to strengthen local features (e.g. PVTv2[33], SRFormer[40], LightViT[10]). Others, from architectural perspective, modified the residual connections by designing inverted residual blocks to enhance cross-layer gradient propagation (e.g. CMT[6]). Partially inspired by the above works, we focus on designing learnable parameters of FFN that interact with input feature maps to dynamically redistribute channel-wise parameters. Combining with SAA, our approach effectively boosts feature extraction capabilities with relatively high computational efficiency.
@@ 2.3 Lightweight ViT
Considering that ViT suffers high computational cost, which hinders its application in resource constrained scenarios. Designing lightweight ViT is a research area of great significance. Current lightweight ViT designs mainly focus on latency reduction on real devices. They optimized GPU and CPU throughputs to reduce the computational overhead of attention mechanisms and overall architecture. RepViT[30] replaced the Transformer’s multi-head self attention (MSA) module with lightweight convolutional blocks. It adopted a multi-branch structure during training to enrich feature representations, and merged these branches into an efficient convolutional layer for inference. Meanwhile, it adjusted the number of channels and network depth to reduce parameters, which realizes lower latency on edge device. MobileViT[21] employed a hybrid architecture with unfold–transform–fold operations, achieving higher accuracy while reducing FLOPs. EfficientViT[1] effectively boosted both accuracy and GPU throughput by revising the inverted bottleneck structure and introducing cascaded group attention.
@@ 3 Method
Our goal is to combine the strength of convolutional neural networks (CNN) with the attention mechanisms, and develop a lightweight ViT based architecture. Following the design principle of pyramid network, our model is divided into 4 stages, each of which produces feature maps of different scales to capture multi-scale image information for detection and segmentation tasks. Between each stage, Patch Embedding layers (which consists of convolution and normalization operations) are adopted to down-sample the resolution to half of its original value. The effectiveness of the convolution stems in improving the accuracies in various vision tasks will be proved in Section 4.5.4. Accordingly, our convolution stem is built upon 2 convolution layers (with kernel size=3, stride=1, padding=1). Meanwhile, we stack various Transformer blocks at each pyramid layer to meet diverse task requirements and computational resource constraints. Each Transformer blocks share the same internal structure, which is comprised of a SAA module and a CIFFN. The overall architecture of our proposed model is shown in Fig. 1.
@@
To construct a hybrid convolution-ViT structure, we connect Local Feature Extraction (LFE)[36] module with modified Transformer block. LFE is an embedding layer with DWSConv and an activation function, which refines the input features to the Transformer block. Inside the Transformer block, SAA module rectifies the sampling strategy by performing the sparse block-wise sampling on the feature map in an adaptive way, which reduces the computational load of traditional attention mechanisms. To deal with information redundancy and limited feature expressivity in high-dimensional features, we employ inter-channel interaction mechanism in FFN. To be specific, a 1×1 convolution is firstly applied to map the input features into a higher dimensional space. Then, depth-wise convolution (DWConv) is used to enhance features within each channel, enriching the feature content. An activation function is then introduced to introduce non-linearity, further improving the model’s fitting ability. During channel interaction, a customized decomposition layer and an element-wise scaling layer are used to redistribute and adjust features across channels, enhancing inter-channel information exchange. Finally, another 1×1 convolution maps the features back to their original dimension. Our approach significantly reduces information redundancy and enhances the model’s capability to represent complex features.
The extracted features are sent to a global average pooling layer to yield global feature representation, followed by a classification layer with a softmax function that ultimately produce the final classification results. Based on this, SAEViT effectively addresses the inefficient feature processing mechanism in traditional ViT, which also provides an efficient and lightweight solution for vision tasks.
@@ 3.1 Sparsely Aggregated Attention
In vision tasks, adjacent image patches often have high redundancies (i.e. they share high similarities). Therefore, we replace the standard attention block with SAA in our lightweight model to mitigate the redundancy and also to capture long-range spatial dependencies, so that the model will be able to filter out and integrate tokens that carry similar information(i.e. reduce the number of tokens and computation costs). SAA is featured by spatial sampling strategy: by down-sampling the tokens with a rate sr, so that only the features in the key regions are retained. To be specific, the stride of down-sampling is set to sr to aggregate the adjacent tokens with high similarity. Then, each non-overlapping sr × sr window is reduced to 1 token through average pooling. It not only reduces the complexity of attention operation, but also preserves the main information within the window through local feature smoothing, avoiding the loss of key details caused by random sampling. This is similar to MaxViT[23]’s "Sparse Global Attention" and Biformer[33]’s "Dynamic Sparse Routing" approach, both of them reduce computational complexity through spatial down-sampling. However, unlike them, SAA later up-samples the feature map via deconvolution so as to recover the original spatial dimension. Deconvolution takes a small sized input feature map and outputs a larger feature map, which restores the spatial structural information through a learnable convolution kernel (i.e. enlarging the size by inserting zeros between input elements and performing standard convolution). For given sizes of input feature maps, if the stride is set to s, then the intermediate expended feature map could be created by inserting s-1 number of zeros between each row and column of each spatial unit. Then, a learnable convolution kernel with a size of k×k is used to convolve the feature map, while the output size is controlled by boundary padding.
@@ 3.2 Channel Interaction Feedforward Network Layer
A standard FFN is typically composed of fully connected layers. However, redundancy exists across different channels when dealing with high-dimensional features. In addition, the linear transformation in fully connected layers applies the same operation to all channels, lacking effective screening mechanism and integration of inter-channel information. As a result, the model may focus on some unimportant information during training while ignoring the correlations between channels, reducing the model’s expressive power and generalization abilities. To address this issue, most existing methods directly implant channel enhancement modules, such as the Squeeze-and-Excitation (SE) module, into the FFN. Meanwhile, the non-linear transformations in conventional FFN often apply activation functions to the outputs of fully connected layers, which may only capture insufficient multi-scale and hierarchical features under complex feature distributions. Under such circumstances, we propose CIFFN, which consists of convolutional layers for channel mapping, DWConv for local information extraction, and a feature decomposition module for decomposing and recombining features. A comparison has been made among FFN, depth-wise separable (DWS) FFN and CIFFN in Fig. 3. The computation flow of CIFFN is expressed as:
@@ 4 Experiments
In this section, multiple visual recognition experiments have been conducted to verify the effectiveness of our proposed model. First, our model is pre-trained on ImageNet-1K[4] for image classification task, which is compared with other popular in terms of accuracy and computational cost. Besides, the performances for object detection and segmentation are tested on COCO [17] and ADE20K[39] respectively. The processing speeds for feature maps are analyzed and compared with other competitors.
@@ 4.1 Image Classification
Results: A comparison of classification results are made between our model and other popular methods, including MobileNetV2[26] and EfficientNet[27] as well as PVT[32], PVTv2[33], and DeiT[28]. All methods are roughly divided into 2 groups based on their number of parameters and floating-point operations (FLOPs). As shown in Table 2, under comparable network size, SAEViT-T and SAEViT-XS achieve Top-1 accuracies of 76.3% and 79.6% on ImageNet-1K with only 0.8 GFLOPs and 1.3 GFLOPs, respectively., which outperform other lightweight models such as MobileNetV2[26], EfficientNet[27], PVT[32], and DeiT-T[28]. Notably, SAEViT-XS surpasses EfficientNet-B1 (79.1% with 0.7 GFLOPs) by 0.5% in accuracy while using marginally more computation, demonstrating its superior efficiency-accuracy trade-off. The performance gain can be attributed to the effective integration of convolutional inductive biases and sparse self-attention, which enhances local feature extraction and reduces redundancy without sacrificing global contextual understanding. Additionally, SAEViT-T realizes 0.5% improvement in accuracy over SBCFormer-XS[19], while SAEViT-XS leads SBCFormer-X[19] by 1.9% in accuracy. Similarly, SAEViT achieves higher precision than Mobile-Former. The reason is that SBCFormer adopts more aggressive spatial downsampling operation to pursue lower CPU latency, while Mobile-Former[2] employs a parallel MobileNet+Transformer architecture with very few tokens to prioritize efficiency. In contrast, our model strikes balanced local-global design under lightweight framework, which makes stronger representational capabilities and Top-1 accuracy improvement.
@@
CaiT-xxs12-withDW[37] 6.6 1.3 75.8
Incepformer-S[22] 16.07 9.41 75.3
@@
CaiT-xxs24-withDW[37] 12.0 2.5 79.6
Incepformer-M[22] 30.96 19.2 77.5